OpenStudy (anonymous):

Given the equation 5x+10= -30, solve for x and choose the correct solution. a. x = -8 b. x = 8 c. x = -4 d. x = 4

OpenStudy (anonymous):

Do you know how to get the 10 over to the other side of the equal sign?

OpenStudy (anonymous):

make a guess at @Dark*Chaos

OpenStudy (anonymous):

nooo...

OpenStudy (anonymous):

rly?

OpenStudy (anonymous):

5x+10= -30 5x +10-10 = -30 -10 subtract 10 from both sides 5x = -40 divide by 5 x = -8

OpenStudy (anonymous):

i will not take a guess i will take a studied estimate lol

OpenStudy (anonymous):

OMG DUDE YOUR A GENIUS

OpenStudy (anonymous):

there's the answer for ya! do you get it though?

OpenStudy (anonymous):

not rly im trying to study it

OpenStudy (anonymous):

OHHH!!! I GET IT!!!

OpenStudy (anonymous):

Openstudy is an excellent place to get answers without explanations =D

OpenStudy (anonymous):

let's try another one... try to solve 3x + 2 = 11

OpenStudy (anonymous):

its basically what you do to one side you do to the other!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

subtract 3 from both sides?

OpenStudy (anonymous):

no...always undo the simpler first

OpenStudy (anonymous):

whoops :( (facepalm) I'm an idiot

OpenStudy (anonymous):

the simpler operations first for this kind of problem...you're not an idiot...just need to practice doing more problems

OpenStudy (anonymous):

for 3x + 2 = 11 undo the addition first, by subtracting from both sides

OpenStudy (anonymous):

omg your like my brother and my teacher at the same time.... my brother is cool and great at math and my teacher makes me nervous xD

OpenStudy (anonymous):

haha... except I'm a girl, but the concept is the same

OpenStudy (anonymous):

WHOOPS SRY!

OpenStudy (anonymous):

did you subtract yet?

OpenStudy (anonymous):

yeah i did

OpenStudy (anonymous):

and...what do we have now?

OpenStudy (anonymous):

can i ask you a weird question?

OpenStudy (anonymous):

i guess, what?

OpenStudy (anonymous):

first of all how old r u and are you a moderator

OpenStudy (anonymous):

No, I'm not a moderator, and I'm 18

OpenStudy (anonymous):

oh ok sweet im 15

OpenStudy (anonymous):

anywho what do boobs feel like?

OpenStudy (anonymous):

... back to the subject: what did you get after you subtracted?

OpenStudy (anonymous):

lol 0_o

OpenStudy (anonymous):

( . y . ) graphic art school here I come!

OpenStudy (anonymous):

lol i got drum roll plz

OpenStudy (anonymous):

1 x = 11

OpenStudy (anonymous):

3x + 2 = 11 subtract 2 from both sides

OpenStudy (anonymous):

whoops 3-2 = 1 and 2- 2 = 0

OpenStudy (anonymous):

don't subtract from 3x, because you already subtracted from the left side...subtract from the right side

OpenStudy (anonymous):

y do u bother to help people on this site?

OpenStudy (anonymous):

because I like to help ppl...IDK why...it's just fun for me

OpenStudy (anonymous):

oh lol thats a good thing.... im sure people appreciate it it like me = )

OpenStudy (anonymous):

anywho was i right =)? btw do u have an fb account or u do and just wont add me lol......

OpenStudy (anonymous):

o_O I guess I can add you... But, yeah, I try to help ppl until they understand on their own. so, subtract 2 from both sides (the left and right side)

OpenStudy (anonymous):

and you are left with 1x = 11

OpenStudy (anonymous):

btw whats your fb name

OpenStudy (anonymous):

no... 3x + 2 = 11 -2 -2 ______________

OpenStudy (anonymous):

solve that

OpenStudy (anonymous):

whoops i subtracted 2 from 3 my bad

OpenStudy (anonymous):

0 and 9

OpenStudy (anonymous):

so now you have : 3x+0 = 9 or 3x=9

OpenStudy (anonymous):

hah she left..... i must seriously be boring.... anywho back to aqw

OpenStudy (anonymous):

oh she didnt leave

OpenStudy (anonymous):

i think i get it now

OpenStudy (anonymous):

unless there is another step

OpenStudy (anonymous):

no...she didn't. 3x=9 undo the multiplication by dividing

OpenStudy (anonymous):

3

OpenStudy (anonymous):

oh yeah and im sry that I asked you what boobs felt like it was very inappropriate on my part....

OpenStudy (anonymous):

IDK..I would think you were just curious

OpenStudy (anonymous):

so, you divided by 3 and got x=3 try this one by yourself: 2x -1 = 7

OpenStudy (anonymous):

hah no i have felt them before but i was just wondering how it felt from a girls perspecive

OpenStudy (anonymous):

2 and 5 so 2x=5

OpenStudy (anonymous):

I would normally just tell you, but IDK if this is a conservative site. I sometimes have weird/curous questions too and it's normal... do you want to try the other problem? 2x-1 =7

OpenStudy (anonymous):

no...look at the number that they are subtracting

OpenStudy (anonymous):

they are subtracting -1 right

OpenStudy (anonymous):

yeah...so you want to do the opposite...you want to add 1 to both sides

OpenStudy (anonymous):

so -1 - -1 would be 0 and 7 - -1 would be 6

OpenStudy (anonymous):

so 0x = 6 and now i divide

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

leave the 2x alone when you subtract

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):